# Constraining anisotropic universe under $f(R,T)$ theory of gravity

### v1

@@ Abstract
We try to find the possibility of a Bianchi V universe in the modified gravitational field theory of $f(R,T)$. We have considered a Lagrangian model in connection between the trace of the energy-momentum tensor $T$ and the Ricci scalar $R$. In order to solve the field equations a power law for the scaling factor was also considered. To make a comparison of the model parameters with the observational data we put constraint on the model under the datasets of the Hubble parameter, Baryon Acoustic Oscillations, Pantheon, joint datasets of Hubble parameter + Pantheon and collective datasets of the Hubble parameter + Baryon Acoustic Oscillations + Pantheon. The outcomes for the Hubble parameter in the present epoch are reasonably acceptable, especially our estimation of this $H_{0}$ is remarkably consistent with various recent Planck Collaboration studies that utilize the $\Lambda$-CDM model.
[…]
[1]
@@ 1 Introduction
In 1998, the discovery of a late-time accelerating cosmos and the subsequent development of dark energy (DE) with repulsive pressure led to a significant advancement in cosmology SupernovaSearchTeam:1998fmf ; SupernovaSearchTeam:2004lze ; SDSS:2003eyi ; Calabrese:2009zza ; Wang:2015tua ; Zhao:2017urm . Numerous cosmological models have been examined to comprehend the characteristics of DE Wang:2016lxa ; Peebles:2002gy ; Padmanabhan:2002ji ; Abdalla:2020ypg ; Li:2012dt . The cosmological constant $\Lambda$ is the most viable choice for dark energy, yet it has certain issues with its theoretical implementation Sahni:2002kh ; Garriga:2000cv ; Frieman:2008sn ; Nojiri:2010wj .Our understanding of dark energy is limited to its phenomenological characteristics, which remain enigmatic and unclear: (i) DE is a cosmic fluid that violates the strong energy condition and has an equation of state parameter. (ii) DE shows a lesser clustering property than dark matter (DM) which is uniformly distributed over the universe on a large cosmological scale. Apart from the $\Lambda$-CDM model, there is also the $X$-CDM model where the dynamical dark energy is parametrized along the spatial direction and the $\phi$-CDM model where dynamical dark energy is represented by a scalar field $\phi$ Samushia:2009ib ; Yashar:2008ju . We note that the $X$-CDM model is confirmed for a considerably larger set of cosmological data, such as growth factor information, BAO distance measurements, and type Ia supernova (SN Ia) apparent magnitude observations. This encourages us to build a Bianchi type I space-time model of the accelerating cosmos.
@@
The $f(R,T)$ theory of gravity, which was put forth by Harko et al. Harko:2011 in 2011, expresses the Lagrangian in terms of the trace of the stress-energy tensor $T$ and the Ricci scalar $R$. It is evident that $f(T)$ and matter function as an effective cosmological constant. Therefore, adding a DE component to the $f(R,T)$ theory of gravity may help explain the Universe’s late time acceleration. Harko Harko:2014pqa has since looked upon a generalized gravity model where the thermodynamical consequences on the matter-geometry coupling will have a definite imprint. The following works are involved with some significant uses of the $f(R,T)$ gravity theory Sharma:2018ikm ; Yadav:2020vih ; Sharma:2019oio ; Sharma:2019hqe . Nojiri and Odintsov Nojiri:2010wj have studied a cosmological model where $f(R)$ is used in place of the action. Carroll et al. Carroll:2003wy observed that the late-time acceleration of our Universe could be represented satisfactorily by cosmological models motivated by $f(R)$ gravity. The authors have looked at feasible cosmological models that meet the requirements of the Solar system test for the $f(R)$ theory of gravity Hu:2007nk . We study an anisotropic singular universe model in $f(R,T)$ gravity with nonminimal matter geometry coupling. The field equations have been precisely solved by accounting for the scale factor’s power law change, which drives $\Lambda$-CDM cosmological scenario. It is important to note that the $f(R,T)$ gravity theory plays an important role in evoking a comprehensive theoretical explanation of the late-time accelerating universe without the assistance of unusual exotic matter or energy.
@@
With the help of the presence of dark energy and dark matter, the $f(R,T)$ theory of gravitation contributes significantly to the explanation of the Universe’s current acceleration. Different models regarding different features of the functional form of this theory have been proposed by several scientists Sharma:2018ikm ; Alvarenga:2013syu ; Sahoo:2017poz . According to this hypothesis, the way that various matter components interact with space-time curvature is a cosmological outcome. Nevertheless, a significant contribution of EMC violation causes rapid growth in the cosmic gravity models of $f(R,T)$. The impact of EMC violation in this theory has not yet been thoroughly investigated. However, a few first investigations on celestial objects by employing this theory have been proposed in the following works Yousaf:2016lls ; Yousaf:2018jkb ; Yousaf:2016awj ; Yousaf:2019zcb ; Das:2016mxq . Inspired by the previous conversations, we now set out to study a cosmological model for the following specifications: (i) there will be a connection between nonminimal matter and geometry, (ii) it will be associated with the space-time of Bianchi V and (iii) here gravity will be defined as $f(R,T)=f_{1}(R)+f_{2}(R)f_{3}(T)$. It is noteworthy that the cosmological implications of a nonminimal coupling under the framework of $f(R,T)$ theory are explored by means of a product between $R$ and $T$ or functions of them. We shall use the simplest coupling $f_{1}(R)=f_{2}(R)=R$ and $f(T)=\zeta T$, with $\zeta$ a constant. Within the $f(R,T)$ gravity formalism, the function $f(R,T)$ has a nontrivial functional form that involves in the nonminimal matter-geometry coupling. Additionally, it benefits from the fact that $\zeta$ = 0 is the retrieval of general relativity. A few practical uses for $f(R,T)=f_{1}(R)+f_{2}(R)f_{3}(T)$ are available in various physical situations Sharma:2018ikm ; Yadav:2019zrw .
@@
The paper is structured as follows: the model and its basic formalism is described in Section 2. The method of observation analysis and observational data sets are given in Section 3. Section 4 deals with the energy conditions and cosmological parameters. Finally, the findings of this paper are summarized in Section 5.
2 Mathematical background of $f(R,T)$ gravity theory
@@
Now, the given equation can be expressed as
@@
whereas the average scale factor can be provided as Sharma et al. Sharma:2019oio
@@
3 Observational Analysis
3.1 Model parameters and observational constraints
@@
1. OHD: From the cosmic chronometric approach, we have collected $55~{}H(z)$ Observed Hubble Data (OHD) points in the interval $0\leq z\leq 2.36$. Table II contains all of the data points from $55~{}H(z)$ of Ref. Lohakare/2022 .
2. BAO: Anisotropic Baryon Acoustic Oscillation (BAO) measurements of $D_{M}(z)/r_{d}$ and $D_{H}(z)/r_{d}$ are included in the final BAO measurements of the SDSS collaboration. Basically this span to eight different redshift intervals (in which $D_{H}(z)=c/H(z)$ is the Hubble distance and $D_{M}(z)$ is the comoving angular diameter distance). Table 3 of Ref. eBOSS:2020yzd compiles all the BAO-only measurements.
3. Pantheon sample: The Pantheon sample was used to get distance modulus measurements of SN Ia Scolnic/2018 . The redshift range $z\in[0.001,2.26]$ contains 1701 light curves that correspond to 1550 distinct SN Ia events.
@@
4 Energy conditions and cosmological Parameters
4.1 Energy Conditions
The energy conditions analogously establish cosmic laws that elucidate the distribution of matter and energy throughout the cosmos, which can be derived from Einstein’s gravitational equations. Thus, we obtain certain conditions, viz. Weak Energy Condition (WEC), Null Energy Condition (NEC), Strong Energy Condition (SEC), and Dominant Energy Condition (DEC) reveal the distribution of matter and energy in the cosmic space.
@@ 4.1 Energy Conditions
    According to the WEC, energy cannot be negative or zero. This requirement keeps the laws of the cosmos impartial and constant. Using the parameters discovered using the Bayesian Analysis, Fig. 14 depicts the WEC’s nature for our model.
@@
    Similar to a more rigid counterpart of the NEC, the SEC guarantees that the energy cannot be negative and sets a boundary for how objects respond to gravity. The cosmos benefits from having this rule in place. Figure 13 depicts the characteristics of SEC for our model utilizing the Bayesian analysis.
@@
    The DEC guarantees that the energy is not just non-negative but also that its distribution cannot be too volatile. It is comparable to asserting that energy cannot go out of control or travel faster than light. Based on the Bayesian Analysis, DEC is shown in Fig. 12.
All the above-mentioned energy conditions combined are as shown in Fig. 15. Figure 8 displays the distribution of energy density $\rho$ wrt time $t$, while Fig. 9 displays the distribution of pressure. Our results demonstrate that, except for SEC the others, e.g., NEC, WEC and DEC are fulfilled. The late-rapid expansion of the universe justifies the violation of SEC. The $f(R,T)$ theory of gravity, having a contribution from the trace energy $T$ can suitably explain the late-time accelerated phase of the universe without invoking any need for the cosmological constant or exotic dark energy as the background agent for the phenomenon.
@@
Figure 10: Graphical presentation of the equation of state parameter vs time.
@@
4.2 State Finder Diagnotics
@@ 4.3 Om(z) parameter
The Hubble parameter is shown by $H_{0}$ in this case. Shahalam et al. Shahalam2015 claim that the negative, zero and positive values of $Om(z)$, respectively, represent the quintessence ($\omega\geq-1$), $\Lambda$-CDM and phantom ($\omega\leq-1$) dark energy theories. However, for the present model, we obtain this parameter as follows:
@@
4.4 Jerk Parameter
After the Hubble and deceleration parameters, the jerk parameter is one that can provide deeper insights into the universe’s history. The universe’s acceleration is tracked throughout time by the jerk parameter, indicated by the symbol $j$. The acceleration is accelerating and speeding up phase when $j$ is positive ($j\geq 0$), whereas the acceleration is said to slow down if $j$ is negative ($j\leq 0$). However, different dark energy theories forecast various jerk parameter values, and thus, scientists can learn more about dark energy by measuring $j$ and comparing it to these predictions. The jerk parameter contributes to the improvement of our models of the cosmos when paired with other cosmic factors. It puts these models to a rigorous examination to ensure they truly reflect what we observe in the physical cosmos. In this context, we note that Fig. 18 depicts the fluctuation of $j$ with respect to redshift $z$ for our model and the derived parameters.
@@ 5 Conclusion
[…]
@@
Overall observation and conclusion can be put forward as follows: the obtained results demonstrate that the proposed model mostly agrees with the observational signatures of the cosmological scenario within a certain range of restrictions.
@@ Acknowledgement
SR would like to express his gratitude to the ICARD facilities at CCASS, GLA University, Mathura, as well as the Visiting Research Associateship Programme at the Inter-University Centre for Astronomy and Astrophysics (IUCAA) in Pune, India.

### v2

@@ Abstract
Abstract: We try to find the possibility of a Bianchi V universe in the modified gravitational field theory of $f(R,T)$. We have considered a Lagrangian model in connection between the trace of the energy-momentum tensor $T$ and the Ricci scalar $R$. In order to solve the field equations a power law for the scaling factor was also considered. To make a comparison of the model parameters with the observational data we put constraint on the model under the datasets of the Hubble parameter, Baryon Acoustic Oscillations, Pantheon, joint datasets of Hubble parameter + Pantheon and collective datasets of the Hubble parameter + Baryon Acoustic Oscillations + Pantheon. The outcomes for the Hubble parameter in the present epoch are reasonably acceptable, especially our estimation of this $H_{0}$ is remarkably consistent with various recent Planck Collaboration studies that utilize the $\Lambda$-CDM model.
@@ I Introduction
In 1998, the discovery of a late-time accelerating cosmos and the subsequent development of dark energy (DE) with repulsive pressure led to a significant advancement in cosmology [1, 2, 3, 4, 5, 6]. Numerous cosmological models have been examined to comprehend the characteristics of DE [7, 8, 9, 10, 11]. The cosmological constant $\Lambda$ is the most viable choice for dark energy, yet it has certain issues with its theoretical implementation [12, 13, 14, 28]. Our understanding of dark energy is limited to its phenomenological characteristics, which remain enigmatic and unclear: (i) DE is a cosmic fluid that violates the strong energy condition and has an equation of state parameter. (ii) DE shows a lesser clustering property than dark matter (DM) which is uniformly distributed over the universe on a large cosmological scale. Apart from the $\Lambda$-CDM model, there is also the $X$-CDM model where the dynamical dark energy is parametrized along the spatial direction and the $\phi$-CDM model where dynamical dark energy is represented by a scalar field $\phi$ [16, 17]. We note that the $X$-CDM model is confirmed for a considerably larger set of cosmological data, such as growth factor information, BAO distance measurements, and type Ia supernova (SN Ia) apparent magnitude observations. This encourages us to build a model of Bianchi type V space-time of the accelerating cosmos.
@@
The $f(R,T)$ theory of gravity, which was put forth by Harko et al.[18] in 2011, expresses the Lagrangian in terms of the trace of the stress-energy tensor $T$ and the Ricci scalar $R$. This paper introduced $f(R,T)$ gravity and explored its applications to cosmology, including explanations for cosmic acceleration and other gravitational phenomena. It is also evident that $f(T)$ and matter function as an effective cosmological constant. Therefore, adding a DE component to the $f(R,T)$ theory of gravity may help explain the Universe’s late time acceleration. Jamil et al. [19] have investigated $f(R,T)$ models to explain accelerated cosmic expansion and presents various cosmological solutions. Later on, the authors of Refs. [20, 21] have described the thermodynamic implications of $f(R,T)$ gravity, focusing on cosmological thermodynamics and the validity of the generalized second law of thermodynamics. Moreover, in 2014, Harko [22] has given a generalised gravity model where the thermodynamical consequences on the matter-geometry coupling will have a definite imprint. It is worthwhile to note that Nojiri et al. [23] have reviewed the various modifications of gravity, including f(R,T) gravity, and examined their applications to late-time cosmic acceleration and early universe inflation. The following works are involved with some significant uses of the $f(R,T)$ gravity theory [24, 25, 26, 27]. Nojiri and Odintsov [28] have studied a cosmological model where $f(R)$ is used in place of the action. Carroll et al. [29] observed that the late time acceleration of our Universe could be represented satisfactorily by cosmological models motivated by $f(R)$ gravity. The authors have looked at feasible cosmological models that meet the requirements of the Solar system test for the $f(R)$ theory of gravity [30]. Therefore, the $f(R,T)$ theory of gravity implies new interactions between geometry and matter and the door is now open to the prospect of investigating matter-geometry coupling, in which the composition of matter may have more nuanced effects on the gravitational field. In this paper, we explore an anisotropic singular universe model in $f(R,T)$ gravity with non-minimal matter geometry coupling. The field equations have been precisely solved by accounting for the scale factor power law change, which drives $\Lambda$-CDM cosmological scenario. Recently Pradhan et al. [31] have reconstructed a model of constant jerk parameter in $f(R,T)$ gravity which explores the late time phenomenon of the universe. Therefore, it is worthwhile to note that the $f(R,T)$ gravity theory plays an important role in evoking a comprehensive theoretical explanation of the late time accelerating universe without an assistance of unusual exotic matter or energy.
@@
With the help of the presence of dark energy and dark matter, the $f(R,T)$ theory of gravitation contributes significantly to the explanation of the Universe’s current acceleration. Different models regarding different features of the functional form of this theory have been proposed by several scientists [24, 32, 33]. According to this hypothesis, the way that various matter components interact with space-time curvature is a cosmological outcome.Nevertheless, a significant contribution of EMC violation causes rapid growth in the cosmic gravity models of $f(R,T)$. The impact of EMC violation in this theory has not yet been thoroughly investigated. However, a few first investigations on celestial objects by employing this theory have been proposed in the following works [34, 35, 36, 37, 38]. Inspired by the previous conversations, we now set out to study a cosmological model for the following specifications: (i) there will be a connection between nonminimal matter and geometry, (ii) it will be associated with the space-time of Bianchi V and (iii) here gravity will be defined as $f(R,T)=f_{1}(R)+f_{2}(R)f_{3}(T)$. It is noteworthy that the cosmological implications of a nonminimal coupling under the framework of $f(R,T)$ theory are explored by means of a product between $R$ and $T$ or functions of them. We shall use the simplest coupling $f_{1}(R)=f_{2}(R)=R$ and $f(T)=\zeta T$, with $\zeta$ a constant. Within the $f(R,T)$ gravity formalism, the function $f(R,T)$ has a nontrivial functional form that involves in the nonminimal matter-geometry coupling. Additionally, it benefits from the fact that $\zeta$ = 0 is the retrieval of general relativity. A few practical uses for $f(R,T)=f_{1}(R)+f_{2}(R)f_{3}(T)$ are available in various physical situations [24, 39].
Among various geometrical models of the universe, Bianchi type V universe represents a natural generalization of the open universe and hence its study is important in the study of fate and dynamics of the universe with non-zero curvature [40, 41]. These models are more general and more adapted to examine certain aspects of the cosmos because it has loosen the assumption of isotropy imposed in case of FLRW model of universe. In 2004, Eriksen et al. [42] have investigated a cosmological phenomenon which violates the statistical isotropy of the universe due to large angle anomalies appear in CMB radiations. In the recent past, several authors have studied Bianchi-V cosmological models in different physical contexts [43, 44, 45, 46]. Furthermore, the studies of the Bianchi–V cosmological models create more interest as these models contain special isotropic cases and permit arbitrarily small levels of anisotropy at some instant of cosmic time which makes these models suitable to describes early phenomenon as well as the late time dynamics of the universe. Therefore, the homogeneous and isotropic FRW cosmological models, which are used to describe standard cosmological models, are a particular case of Bianchi–V universe, depending on whether the negatively curved spatial geometry and possibility of forming distortion due to cosmic shear. Recently Goswami et al. [47] have investigated a bulk viscous Bianchi type V universe and constrained its model parameters with recent H(z) and Pantheon compilation of SN Ia observational data. Some important applications of cosmological models in Bianchi V space - time are given in Refs. [48, 49, 50, 51] specially in Ref [51], the authors have examined the effects of anisotropic shear and implications of Bianchi type V models with cosmic microwave background (CMB) anisotropies. Moreover, Hewitt et al. [52] have investigated a dynamical system which approaches anisotropic Bianchi V cosmology and explored the phase space of Bianchi type V models to search its long-term behavior.
@@
The paper is structured as follows: the model and its basic formalism is described in Section II. The method of observation analysis and observational data sets are given in Section III. Section IV deals with the energy conditions and cosmological parameters. Finally, the findings of this paper are summarized in Section V.
II Mathematical background of $f(R,T)$ gravity theory
@@
Now the given equation can be expressed as
@@
whereas the average scale factor can be provided as Sharma et al. [26]
@@
III Observational Analysis
III.1 Model parameters and observational constraints
@@
1. OHD: From the cosmic chronometric (CC) and BAO approach, we have collected $57~{}H(z)$ Observed Hubble Data (OHD) points in the interval $0\leq z\leq 2.36$. Table II contains $55$ $H(z)$ data points [53] while all $57~{}H(z)$ observational data points with corresponding errors from CC (31 points), and BAO (26 points) are given in Ref. [54].
2. BAO: Anisotropic Baryon Acoustic Oscillation (BAO) measurements of $D_{M}(z)/r_{d}$ and $D_{H}(z)/r_{d}$ are included in the final BAO measurements of the SDSS collaboration. Basically this span to eight different redshift intervals (in which $D_{H}(z)=c/H(z)$ is the Hubble distance and $D_{M}(z)$ is the comoving angular diameter distance). Table 3 of Ref. [55] compiles all the BAO-only measurements.
3. Pantheon sample: The Pantheon sample was used to get distance modulus measurements of SN Ia [56]. The redshift range $z\in[0.001,2.26]$ contains 1701 light curves that correspond to 1550 distinct SN Ia events.
@@
IV Energy conditions and cosmological Parameters
IV.1 Energy Conditions
The energy conditions, analogously establish cosmic laws that elucidate the distribution of the matter and energy throughout the cosmos which can be derived from Einstein’s gravitational equations. Thus we obtain certain conditions, viz. Weak Energy Condition (WEC), Null Energy Condition (NEC), Strong Energy Condition (SEC) and Dominant Energy Condition (DEC) which reveal the distribution of matter and energy in the cosmic space.
@@ IV.1 Energy Conditions
    According to the WEC, energy can nowhere be negative or zero. This requirement keeps the laws of the cosmos impartial and constant. Using the parameters discovered using the Bayesian Analysis, Fig. 14 depicts the WEC’s nature for our model.
@@
    Similar to a more rigid counterpart of the NEC, the SEC which guarantees that the energy cannot be negative and sets a boundary for how objects respond to gravity. The cosmos benefits from having this rule in place. Figure 13 depicts the characteristics of SEC for our model utilizing the Bayesian analysis.
@@
    The DEC guarantees that the energy is not just non-negative but also that its distribution cannot be too volatile. It is comparable to asserting that the energy cannot go out of control or travel faster than light. Based on the Bayesian Analysis DEC is shown in Fig. 12.
All the above-mentioned energy conditions combinedly are as shown in the Fig. 15. Figure 8 displays the distribution of energy density $\rho$ versus time $t$, while Fig. 9 displays the distribution of pressure. Our results demonstrate that, except for SEC the others, e.g. NEC, WEC and DEC are fulfilled. The late time rapid expansion of the universe justifies violation of SEC. The $f(R,T)$ theory of gravity having contribution from the trace energy $T$ can suitably explain the late time accelerated phase of the universe without invoking any need for the cosmological constant or exotic dark energy as the background agent for the phenomenon.
@@
Figure 10: Graphical presentation of equation of state parameter vs time.
@@
IV.2 State Finder Diagnotics
@@ IV.3 Om(z) parameter
The Hubble parameter is shown by $H_{0}$ in this case. Shahalam et al. [57] claim that the negative, zero and positive values of $Om(z)$, respectively, represent the quintessence ($\omega\geq-1$), $\Lambda$-CDM and phantom ($\omega\leq-1$) dark energy theories. However, for the present model, we obtain this parameter as follows:
@@
IV.4 Jerk Parameter
After the Hubble and deceleration parameters, the jerk parameter is one which can provide deeper insights into the universe’s history. The universe’s acceleration is tracked throughout time by the jerk parameter, indicated by the symbol $j$. The acceleration is accelerating and speeding up phase when $j$ is positive ($j\geq 0$) whereas the acceleration is said to slow down if $j$ is negative ($j\leq 0$). However, different dark energy theories forecast various jerk parameter values and thus scientists can learn more about dark energy by measuring $j$ and comparing it to these predictions. The jerk parameter contributes to the improvement of our models of the cosmos when paired with other cosmic factors. It puts these models to a rigorous examination to ensure they truly reflect what we observe in the physical cosmos. In this context we note that Fig. 18 depicts the fluctuation of $j$ with respect to redshift $z$ for our model and the derived parameters.
@@ V Conclusion
(i) To compare the model parameters with the observational data we impose constraint on the model using the data sets from the Hubble parameter $H(z)$, BAO, Pantheon, combined $H(z)$ + Pantheon, and combined $H(z)$ + BAO + Pantheon. The outcomes for $H_{0}$ under our $f(R,T)$ gravity model are as follows: $H_{0}=68.790^{+2.250}_{-2.046}$ km s-1 Mpc-1, $H_{0}=69.499^{+2.561}_{-1.863}$ km s-1 Mpc-1, $H_{0}=69.102^{+2.453}_{-2.072}$ km s-1 Mpc-1, $H_{0}=67.633^{+2.448}_{-2.013}$ km s-1 Mpc-1, $H_{0}=68.381^{+2.219}_{-2.032}$ km s-1 Mpc-1, respectively.
@@
Overall statistical analysis and conclusion can be put forward as follows: the obtained results demonstrate that the proposed model mostly agrees with the observational signatures of the cosmological scenario within a certain range of restrictions.
@@ Acknowledgement
We are very much grateful to the honorable referee and to the editor for the illuminating suggestions that have significantly improved our work in terms of research quality and presentation. SR would like to express his gratitude to the ICARD facilities at CCASS, GLA University, Mathura, as well as the Visiting Research Associateship Programme at the Inter-University Centre for Astronomy and Astrophysics (IUCAA) in Pune, India. AKY is grateful to Dr. Nafis Ahmad, King Khalid University for fruitful discussions. The author (AKY) extends his appreciation to the Deanship of Scientific Research at King Khalid University, Saudi Arabia for funding this work through large group Research Project under grant number RGP. 2/315/45.